OpenStudy (anonymous):

If f(x)=1/x+2 What is f(3+h)? a) (3+h)/(x+2) b) 1/ (x+2) (3+h) c) 1/ 3+h d) 1/ h+5

OpenStudy (nincompoop):

what is h?

OpenStudy (shamil98):

Doesn't matter, just plug it in.

OpenStudy (nincompoop):

plug things in

OpenStudy (shamil98):

\[\huge f(3+h) = \frac{ 1 }{ (3+h)+2 }\] Simplify it.

OpenStudy (nincompoop):

1/x+2 means (1/x) + 2

OpenStudy (shamil98):

I think he meant it to be: \[\huge f(x) = \frac {1} {x+2}\]

OpenStudy (nincompoop):

laughing out loud

OpenStudy (shamil98):

but next time use 1/(x+2) to clarify it..

OpenStudy (anonymous):

I think given that, the right answer would be a? (3+h)/(x+2)

OpenStudy (anonymous):

what do you get when you simplify \[\frac{ 1 }{ (3+h)+2 }\]